# Mind the (DH) Gap! A Contrast in Risky Choices Between Reasoning and Conversational LLMs

### v1

@@ C.4 Additional Details on Impact of Training and Scale for Open Models
Here, we present the complete results for the sequential training checkpoints—SFT, DPO, and RLVF (final)—for both the instruction and thinking variants of the open Olmo3 models. Their training details can be found in the technical report Ettinger et al. (2025). The results, reported in Tables 5 and 6, do not indicate any systematic variation attributable to either the DPO or RLVF training stages. Again, given that Olmo3-7B-Instruct-SFT is initialized from Olmo3-7B-Think-SFT, the observed parameter differences are consistent with the hypothesis that the initial SFT stage may be the primary source of divergence between the two model variants.
Olmo-3-7B-Instruct-SFT 0.89 0.14 0.81 318 0.81 0.024
Olmo-3-7B-Instruct-DPO 0.90 0.13 0.81 301 0.80 0.032
Olmo-3-7B-Instruct 0.87 0.20 0.82 220 0.73 0.042
Olmo-3-7B-Think-SFT 1.04 0.51 1.05 1000 1.00 0.001
Olmo-3-7B-Think-DPO 0.87 1000 1.16 1000 1.00 0.000
Olmo-3-7B-Think 1.33 1000 1.15 1000 1.00 0.001
Olmo-3-32B-Think-SFT 1.33 1000 1.16 1000 1.00 0.000
Olmo-3-32B-Think-DPO 1.32 1000 1.15 1000 1.00 0.000
Olmo-3-32B-Think 1.33 1000 1.16 1000 1.00 0.000
@@
Table 5: Prospect theory model parameters for models over sequential training stages (explicit prospects).
Olmo-3-7B-Instruct-SFT 166 0.01 1.04 0.70 0.39 0.020
Olmo-3-7B-Instruct-DPO 188 0.17 1.04 0.68 0.39 0.019
Olmo-3-7B-Instruct 1.25 0.27 1.00 4.80 0.44 0.020
Olmo-3-7B-Think-SFT 0.75 0.50 1.08 20.8 0.69 0.097
Olmo-3-7B-Think-DPO 0.93 3.65 0.83 1000 0.98 0.010
Olmo-3-7B-Think 0.82 0.69 0.78 702 1.00 0.001
Olmo-3-32B-Think-SFT 0.95 0.78 0.87 1000 0.95 0.023
Olmo-3-32B-Think-DPO 0.95 0.64 0.88 1000 0.99 0.003
Olmo-3-32B-Think 0.76 1000 1.31 405 0.95 0.023
@@
Table 6: Prospect theory model parameters for models over sequential training stages (implicit prospects).
C.5 Ablation and Alternative Model Results
@@
This section presents the empirical results for the restricted and alternative behavioral models defined in Appendix B.2.3. We evaluate the impact of isolating decisiveness versus risk preferences by comparing the Beta-Only (Model 1) and Shape-Only (Model 2) variants. Additionally, we evaluate Regret Theory (Model 4) as an alternative mechanism to Prospect Theory.
C.5.1 Parameter Estimates: Restricted Prospect Theory Models
@@
Table 7 and 8 details the fitted parameters for the prospect theory ablation models. For Model 1, we report only $\beta$ (fixing $\sigma=\lambda=\gamma=1$). For Model 2, we report the shape parameters (fixing $\beta=1$).
@@
C.5.2 Parameter Estimates: Regret Aversion
Table 9 and 10 presents the fitted parameters for Model 4, the regret aversion model. Here, $\lambda_{reg}$ represents decisiveness in the regret framework, while $\kappa$ and $\alpha$ control the shape of the regret function.
@@ C.5.2 Parameter Estimates: Regret Aversion
C.6 Confidence Intervals
Tables 111213, and 14 presents the 95% confidence intervals for the parameters of the four-parameter prospect theory model (Model 3), estimated via the parametric bootstrap method as in Appendix B.2.3 with $B=1000$ iterations.
Model $\sigma$ (95% CI) $\lambda$ (95% CI) $\gamma$ (95% CI) $\beta$ (95% CI)
DeepSeek-R1 [0.79, 1.26] [0.83, 1000] [0.81, 1.13] [648, 1000]
Gemini-2.5-Pro [0.77, 0.94] [0.34, 2.57] [0.82, 0.97] [1000, 1000]
GPT-5.1 [1.23, 1.36] [1000, 1000] [1.13, 1.23] [1000, 1000]
Claude-4.5-Haiku [0.89, 1.36] [1.63, 27] [0.82, 1.33] [16, 188]
DeepSeek-Chat [0.79, 0.87] [0.21, 0.85] [0.82, 0.88] [159, 159]
Gemini-2.5-Flash [0.76, 1.56] [0.10, 9] [0.81, 2.79] [6, 168]
GPT-4.1 [0.50, 0.70] [0.01, 0.10] [0.97, 2.00] [150, 150]
Human [0.01, 3.81] [0.01, 1000] [0.03, 1.96] [0.02, 194]
@@ C.6 Confidence Intervals
Table 11: 95% Confidence Intervals for prospect theory parameters of black-box models with explicit prospects.
Model $\sigma$ (95% CI) $\lambda$ (95% CI) $\gamma$ (95% CI) $\beta$ (95% CI)
DeepSeek-R1 [1.05, 1.06] [0.83, 2.34] [0.91, 0.94] [1000, 1000]
Gemini-2.5-Pro [0.83, 0.86] [0.81, 1.40] [1.20, 1.28] [128, 128]
GPT-5.1 [1.04, 1.06] [1000, 1000] [0.92, 0.94] [1000, 1000]
Claude-4.5-Haiku [0.62, 0.80] [0.64, 1.27] [1.95, 2.28] [9, 16]
DeepSeek-Chat [0.62, 1.06] [0.27, 0.87] [2.19, 3.03] [6, 13]
Gemini-2.5-Flash [1.24, 1.39] [0.65, 1.28] [0.65, 0.70] [13, 21]
GPT-4.1 [1.09, 1.16] [2.46, 4.90] [0.62, 0.67] [14, 22]
Human [0.74, 1.24] [0.21, 2.81] [0.71, 1.44] [3.32, 16]
@@
Table 12: 95% Confidence Intervals for prospect theory parameters of black-box models with implicit prospects.
Model $\sigma$ (95% CI) $\lambda$ (95% CI) $\gamma$ (95% CI) $\beta$ (95% CI)
Qwen2.5-7B-Instruct [0.09, 1.39] [0.04, 20] [1.97, 25] [1.51, 117]
Qwen2.5-Math-7B [0.85, 1.26] [0.21, 0.61] [0.82, 1.22] [1000, 1000]
Qwen3-30B-Instruct [0.83, 2.33] [0.01, 0.29] [0.79, 3.09] [8, 197]
Qwen3-30B-Thinking [0.79, 1.33] [1000, 1000] [0.80, 1.16] [1000, 1000]
Olmo-3-7B-Instruct-SFT [0.83, 0.89] [0.05, 0.28] [0.80, 0.83] [318, 318]
Olmo-3-7B-Instruct-DPO [0.82, 0.89] [0.04, 0.28] [0.80, 0.85] [301, 301]
Olmo-3-7B-Instruct [0.83, 0.89] [0.07, 0.39] [0.80, 0.84] [220, 220]
Olmo-3-7B-Think-SFT [1.07, 1.43] [0.36, 1.17] [1.02, 1.32] [1000, 1000]
Olmo-3-7B-Think-DPO [1.23, 1.36] [1000, 1000] [1.13, 1.23] [1000, 1000]
Olmo-3-7B-Think [1.23, 1.35] [1000, 1000] [1.13, 1.23] [1000, 1000]
Olmo-3-32B-Think-SFT [1.23, 1.36] [1000, 1000] [1.13, 1.24] [1000, 1000]
Olmo-3-32B-Think-DPO [1.23, 1.35] [1000, 1000] [1.12, 1.27] [1000, 1000]
Olmo-3-32B-Think [1.23, 1.35] [1000, 1000] [1.12, 1.23] [1000, 1000]
@@
Table 13: 95% Confidence Intervals for prospect theory parameters of open-weight models with explicit prospects.
Model $\sigma$ (95% CI) $\lambda$ (95% CI) $\gamma$ (95% CI) $\beta$ (95% CI)
Qwen2.5-7B-Instruct [1.31, 2.06] [0.12, 0.32] [2.78, 3.79] [14, 27]
Qwen2.5-Math-7B [0.15, 1.84] [0.01, 0.41] [0.73, 1.61] [4.23, 11]
Qwen3-30B-Instruct [1.00, 1.23] [0.04, 0.36] [0.76, 0.89] [13, 22]
Qwen3-30B-Thinking [0.97, 0.98] [0.80, 3.29] [0.85, 0.86] [914, 914]
Olmo-3-7B-Instruct-SFT [263, 263] [0.01, 0.42] [0.29, 5] [0.46, 3.05]
Olmo-3-7B-Instruct-DPO [309, 309] [0.01, 0.19] [2.96, 18] [0.90, 21]
Olmo-3-7B-Instruct [1.00, 1.84] [0.01, 0.72] [0.75, 2.15] [2.45, 8]
Olmo-3-7B-Think-SFT [0.64, 0.95] [0.26, 0.97] [0.86, 1.40] [15, 36]
Olmo-3-7B-Think-DPO [0.93, 1.00] [3.65, 8] [0.82, 0.87] [1000, 1000]
Olmo-3-7B-Think [0.43, 1.37] [0.31, 3.18] [0.53, 1.40] [702, 702]
Olmo-3-32B-Think-SFT [0.90, 1.04] [0.41, 2.28] [0.86, 0.96] [1000, 1000]
Olmo-3-32B-Think-DPO [0.89, 1.38] [0.38, 1.60] [0.87, 1.49] [1000, 1000]
Olmo-3-32B-Think [0.76, 0.92] [1000, 1000] [1.15, 1.64] [406, 406]
@@
Table 14: 95% Confidence Intervals for prospect theory parameters of open-weight models with implicit prospects.

### v2

@@ B.3 Prompts and Prompts Sensitivity Tests
Additionally, we conducted prompt sensitivity analysis using three variations: "You will be provided with…", "Here are…", "Please consider…". We collected data from 4 RMs (gpt-5.1, gemini-2.5-pro, Olmo3-7B-think,Qwen2.5-Math-instruct) and 4 CMs (gpt-4.1, gemini-2.5-flash, Olmo3-7B-instruct,Qwen2.5-instruct). Overall, we found that the RM-CM gap is robust across prompt variations. In particular, the range of RM–RM correlation is (0.82 to 0.99) with mean 0.91, whereas the range of RM–CM correlation is (0.21 to 0.78) with mean of 0.31.
@@ C.4 Additional Details on Impact of Training and Scale for Open Models
Here, we present the complete results for the sequential training checkpoints—SFT, DPO, and RLVF (final)—for both the instruction and thinking variants of the open Olmo3 models. Their training details can be found in the technical report Ettinger et al. (2025). The results for the dual-beta model, reported in Tables 5 and 6, do not indicate any systematic variation attributable to either the DPO or RLVF training stages. Again, given that Olmo3-7B-Instruct-SFT is initialized from Olmo3-7B-Think-SFT, the observed parameter differences are consistent with the hypothesis that the initial SFT stage may be the primary source of divergence between the two model variants.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
Olmo-3-7B-Instruct-SFT 0.88 0.80 309 130 0.80 0.029
Olmo-3-7B-Instruct-DPO 0.88 0.81 320 142 0.82 0.030
Olmo-3-7B-Instruct 0.88 0.81 215 121 0.73 0.041
Olmo-3-7B-Think-SFT 1.27 1.13 1000 509 1.00 0.001
Olmo-3-7B-Think-DPO 1.33 1.16 1000 1000 1.00 0.000
Olmo-3-7B-Think 1.33 1.16 1000 1000 1.00 0.001
Olmo-3-32B-Think-SFT 1.33 1.16 1000 1000 1.00 0.000
Olmo-3-32B-Think-DPO 1.33 1.16 1000 1000 1.00 0.000
Olmo-3-32B-Think 1.33 1.16 1000 1000 1.00 0.000
@@
Table 5: Dual-beta prospect theory model parameters for models over sequential training stages (explicit prospects).
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$ Corr MSE
Olmo-3-7B-Instruct-SFT 2.40 1.26 1.41 0.01 0.35 0.021
Olmo-3-7B-Instruct-DPO 1.50 1.05 2.86 0.07 0.37 0.020
Olmo-3-7B-Instruct 1.25 0.99 4.80 1.29 0.44 0.020
Olmo-3-7B-Think-SFT 0.78 1.04 21.7 10.6 0.61 0.097
Olmo-3-7B-Think-DPO 0.98 0.86 1000 824 0.96 0.015
Olmo-3-7B-Think 0.73 0.71 1000 466 1.00 0.001
Olmo-3-32B-Think-SFT 0.88 0.95 418 355 0.95 0.023
Olmo-3-32B-Think-DPO 0.96 0.89 1000 563 0.99 0.003
Olmo-3-32B-Think 0.95 1.07 1000 235 0.93 0.032
@@
Table 6: Dual-beta prospect theory model parameters for models over sequential training stages (implicit prospects).
C.5 Confidence Intervals
Tables 789, and 10 present the 95% confidence intervals for the dual-beta prospect theory model, estimated via the parametric bootstrap method described in Appendix B.2.3 with $B=1000$ iterations.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$
DeepSeek-R1 [0.81, 0.96] [0.80, 0.91] [832, 832] [711, 711]
Gemini-2.5-Pro [0.76, 0.86] [0.81, 0.88] [951, 951] [1000, 1000]
GPT-5.1 [1.28, 1.35] [1.14, 1.22] [1000, 1000] [1000, 1000]
Claude-4.5-Haiku [0.88, 0.91] [0.79, 0.81] [221, 221] [244, 244]
DeepSeek-Chat [0.84, 0.89] [0.81, 0.84] [181, 181] [149, 149]
Gemini-2.5-Flash [0.84, 3.27] [0.81, 4.61] [4.53, 303] [15, 1000]
GPT-4.1 [0.52, 0.95] [1.33, 2.48] [202, 282] [0.01, 42]
Human [0.32, 3.59] [0.01, 2.21] [7, 379] [0.01, 189]
Table 7: 95% Confidence Intervals for dual-beta prospect theory parameters of black-box models with explicit prospects.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$
DeepSeek-R1 [0.97, 1.01] [0.98, 1.04] [1000, 1000] [346, 346]
Gemini-2.5-Pro [0.96, 0.99] [0.97, 1.02] [147, 147] [100, 100]
GPT-5.1 [0.99, 1.03] [0.96, 1.02] [460, 460] [1000, 1000]
Claude-4.5-Haiku [1.01, 1.14] [0.92, 1.10] [3.27, 6.59] [19.1, 28.1]
DeepSeek-Chat [0.68, 0.96] [0.68, 0.78] [8.43, 14.0] [0.01, 1.24]
Gemini-2.5-Flash [1.46, 1.68] [0.77, 0.89] [1.29, 3.89] [11.8, 17.2]
GPT-4.1 [1.44, 1.59] [0.72, 0.77] [0.01, 1.57] [34.6, 47.5]
Human [0.78, 1.19] [0.79, 1.77] [3.68, 11.1] [3.71, 11.7]
Table 8: 95% Confidence Intervals for dual-beta prospect theory parameters of black-box models with implicit prospects.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$
Qwen2.5-7B-Instruct [0.30, 0.90] [0.35, 0.78] [4.86, 140] [0.01, 16.2]
Qwen2.5-Math-7B [0.80, 0.96] [0.81, 0.98] [1000, 1000] [403, 403]
Qwen3-30B-Instruct [0.88, 2.06] [0.81, 3.40] [15.9, 213] [0.01, 17.2]
Qwen3-30B-Thinking [0.82, 0.92] [0.80, 0.90] [726, 726] [1000, 1000]
Olmo-3-7B-Instruct-SFT [0.86, 0.89] [0.80, 0.81] [309, 309] [130, 130]
Olmo-3-7B-Instruct-DPO [0.87, 0.89] [0.80, 0.82] [320, 320] [142, 142]
Olmo-3-7B-Instruct [0.86, 0.89] [0.80, 0.82] [215, 215] [121, 121]
Olmo-3-7B-Think-SFT [0.93, 1.71] [0.85, 1.58] [1000, 1000] [509, 509]
Olmo-3-7B-Think-DPO [1.28, 1.35] [1.14, 1.22] [1000, 1000] [1000, 1000]
Olmo-3-7B-Think [1.28, 1.35] [1.13, 1.22] [1000, 1000] [1000, 1000]
Olmo-3-32B-Think-SFT [1.28, 1.35] [1.13, 1.22] [1000, 1000] [1000, 1000]
Olmo-3-32B-Think-DPO [1.27, 1.34] [1.14, 1.23] [1000, 1000] [1000, 1000]
Olmo-3-32B-Think [1.28, 1.35] [1.14, 1.22] [1000, 1000] [1000, 1000]
Table 9: 95% Confidence Intervals for dual-beta prospect theory parameters of open-weight models with explicit prospects.
Model $\boldsymbol{\sigma}$ $\boldsymbol{\gamma}$ $\boldsymbol{\beta}_{\text{gain}}$ $\boldsymbol{\beta}_{\text{loss}}$
Qwen2.5-7B-Instruct [23.4, 198] [0.47, 0.90] [0.01, 0.29] [2.07, 3.28]
Qwen2.5-Math-7B [0.20, 0.77] [0.76, 1.48] [4.77, 9.88] [0.01, 2.02]
Qwen3-30B-Instruct [1.02, 1.18] [0.78, 0.89] [13.8, 20.2] [1.04, 4.84]
Qwen3-30B-Thinking [0.99, 1.00] [0.87, 0.88] [668, 668] [585, 585]
Olmo-3-7B-Instruct-SFT [1.45, 54.5] [0.63, 5.88] [0.62, 3.47] [0.01, 0.76]
Olmo-3-7B-Instruct-DPO [1.19, 4.52] [0.75, 3.52] [1.40, 4.91] [0.01, 1.26]
Olmo-3-7B-Instruct [1.08, 1.66] [0.80, 2.07] [3.10, 7.07] [0.01, 2.82]
Olmo-3-7B-Think-SFT [0.68, 0.92] [0.89, 1.61] [16.5, 30.5] [8.02, 16.0]
Olmo-3-7B-Think-DPO [0.97, 1.18] [0.85, 0.99] [1000, 1000] [824, 824]
Olmo-3-7B-Think [0.47, 1.11] [0.55, 0.97] [1000, 1000] [466, 466]
Olmo-3-32B-Think-SFT [0.78, 0.92] [0.92, 1.12] [418, 418] [355, 355]
Olmo-3-32B-Think-DPO [0.93, 1.42] [0.88, 1.45] [1000, 1000] [563, 563]
Olmo-3-32B-Think [0.91, 1.00] [1.00, 1.39] [1000, 1000] [235, 235]
Table 10: 95% Confidence Intervals for dual-beta prospect theory parameters of open-weight models with implicit prospects.
C.6 Ablation and Alternative Model Results
@@
This section presents results for alternative behavioral models defined in Appendix B.2.3: two restricted variants of the standard PT parameterization ($\sigma,\lambda,\gamma,\beta$) and a regret aversion model.
C.6.1 Parameter Estimates: Restricted Prospect Theory Models
@@
Table 11 and 12 detail the fitted parameters for the prospect theory ablation models. For Model 1, we report only $\beta$ (fixing $\sigma=\lambda=\gamma=1$). For Model 2, we report the shape parameters (fixing $\beta=1$).
@@
C.6.2 Parameter Estimates: Regret Aversion
Table 13 and 14 present the fitted parameters for Model 4, the regret aversion model. Here, $\lambda_{reg}$ represents decisiveness in the regret framework, while $\kappa$ and $\alpha$ control the shape of the regret function.
@@ C.6.2 Parameter Estimates: Regret Aversion
C.7 Standard Prospect Theory Results
As discussed in Appendix B.2.1, we also fit a standard prospect theory parameterization with four parameters: $\sigma$ (risk preference), $\lambda$ (loss aversion), $\gamma$ (probability weighting), and $\beta$ (decisiveness). While this formulation is widely used, $\lambda$ and $\beta$ are not jointly identifiable in pure-loss prospects, as the model can only recover their product (see Section B.2.1 for details). We report these results here for completeness; qualitative conclusions are consistent with the dual-beta specification reported in the main text.
C.7.1 Frontier Models
DeepSeek-R1 0.79 1000 0.86 1000 0.98 0.012
Gemini-Pro 0.82 0.57 0.83 1000 0.95 0.019
GPT-5.1 1.33 1000 1.16 1000 1.00 0.002
Claude Haiku 0.92 3.89 0.83 89 0.72 0.063
DeepSeek-Chat 0.84 0.43 0.84 159 0.75 0.042
Gemini-Flash 0.95 1.58 0.94 20 0.87 0.015
GPT-4.1 0.51 0.01 1.30 150 0.53 0.053
Human 0.04 66 0.13 0.49 0.67 0.004
@@ C.7.1 Frontier Models
Table 15: Standard prospect theory parameters for frontier models with explicit prospects.
DeepSeek-R1 1.05 1.20 0.92 1000 0.99 0.005
Gemini-Pro 0.84 1.05 1.24 128 0.95 0.021
GPT-5.1 1.04 1000 0.94 1000 1.00 0.001
Claude Haiku 0.70 0.90 2.10 12 0.66 0.059
DeepSeek-Chat 0.80 0.48 2.51 9 0.59 0.039
Gemini-Flash 1.31 0.91 0.67 17 0.69 0.037
GPT-4.1 1.12 3.42 0.64 18 0.69 0.080
Human 1.03 0.89 0.83 9 0.55 0.022
@@
Table 16: Standard prospect theory parameters for frontier models with implicit prospects.
C.7.2 Open Models
Qwen2.5-7B-Instruct 0.58 0.80 3.59 8 0.83 0.017
Qwen2.5-Math-7B 1.11 0.34 1.03 1000 0.99 0.004
Qwen3-30B-Instruct 0.94 0.01 0.86 48 0.57 0.041
Qwen3-30B-Thinking 0.87 1000 0.81 1000 0.99 0.003
Olmo-3-7B-Instruct 0.87 0.20 0.82 220 0.73 0.042
Olmo-3-7B-Think 1.33 1000 1.15 1000 1.00 0.001
Olmo-3-32B-Think 1.33 1000 1.16 1000 1.00 0.000
@@ C.7.2 Open Models
Table 17: Standard prospect theory parameters for open models with explicit prospects.
Qwen2.5-7B-Instruct 1.60 0.21 3.17 19 0.58 0.056
Qwen2.5-Math-7B 0.55 0.12 0.88 7 0.54 0.027
Qwen3-30B-Instruct 1.11 0.19 0.81 17 0.50 0.047
Qwen3-30B-Thinking 0.98 1.67 0.85 913 0.95 0.024
Olmo-3-7B-Instruct 1.25 0.27 1.00 4.80 0.44 0.020
Olmo-3-7B-Think 0.82 0.69 0.78 702 1.00 0.001
Olmo-3-32B-Think 0.76 1000 1.31 405 0.95 0.023
@@
Table 18: Standard prospect theory parameters for open models with implicit prospects.